### v1

# When is Offline Policy Selection Sample Efficient for Reinforcement Learning?

Abstract
Offline reinforcement learning algorithms often require careful hyperparameter tuning. Consequently, before deployment, we need to select amongst a set of candidate policies. As yet, however, there is little understanding about the fundamental limits of this offline policy selection (OPS) problem. In this work we aim to provide clarity on when sample efficient OPS is possible, primarily by connecting OPS to off-policy policy evaluation (OPE) and Bellman error (BE) estimation. We first show a hardness result, that in the worst case, OPS is just as hard as OPE, by proving a reduction of OPE to OPS. As a result, no OPS method can be more sample efficient than OPE in the worst case. We then propose a BE method for OPS, called Identifiable BE Selection (IBES), that has a straightforward method for selecting its own hyperparameters. We highlight that using IBES for OPS generally has more requirements than OPE methods, but if satisfied, can be more sample efficient. We conclude with an empirical study comparing OPE and IBES, and by showing the difficulty of OPS on an offline Atari benchmark dataset.
Keywords: Reinforcement learning, offline reinforcement learning, offline policy selection, off-policy policy evaluation
1 Introduction
Offline reinforcement learning (RL)—learning a policy from a dataset—is useful for many real-world applications, where learning from online interaction may be expensive or dangerous (Levine et al., 2020). There have been significant advances in offline policy learning algorithms with demonstrations that performant policies can be learned purely from offline data (Agarwal et al., 2020). Successful use of these algorithms, however, requires careful hyperparameter selection (Wu et al., 2019; Gulcehre et al., 2020; Kumar et al., 2021).
Despite the fact that it is essential, hyperparameter selection for offline policy learning algorithms remains relatively unexplored. One of the reasons is that it is inherently hard, as we formalize in this paper. Unlike supervised learning where we can use validation performance, in offline RL it is hard to get an accurate measure of policy performance. In fact, it is well known that off-policy policy evaluation (OPE)—estimating the performance of a policy from offline data—is hard (Wang et al., 2021). To emphasize this fact, in Figure 1, we visualize the poor correlation between the true performance and performance estimates using a standard OPE approach called Fitted Q Evaluation (FQE) (Le et al., 2019) . The datasets are not adversarially designed; rather they are two Atari datasets that were designed to provide reasonable data for offline learning algorithms.
Figure 1: Correlation between true performance and estimated performance on two Atari datasets. We learn 90 policies offline using an algorithm called CQL (Kumar et al., 2020) and evaluate these policies using FQE on 5 different random datasets. These policies are generated by running CQL with a variety of different choices for two hyperparameters: the number of training steps and the conservative parameter. Each point in the scatter plot corresponds to a (policy, evaluation dataset) pair. The x-axis is the actual policy performance for that policy and y-axis is the estimated policy performance for that policy using that evaluation dataset. Colors represent different random seeds, namely different evaluation datasets. The Kendall rank correlation coefficient is shown in the label of the plot. If the FQE estimates accurately rank policies, we expect to see a strong linear relationship and a Kendall rank correlation coefficient close to $1$. Neither of these behaviors are seen here, and it is clear FQE does not provide an effective mechanism to rank policies.
Consequently, most offline RL papers sidestep the issue of hyperparameter selection and show results for ideal performance. In other words, they tune their hyperparameters by checking the performance of the policy in the environment. Consider, for example, the Conservative Q-learning (CQL) algorithm (Kumar et al., 2020), used to obtain the policies for Figure 1. For their experiments, CQL is built on top of already-tuned online algorithms for their environments, which implicitly relies on tuning using the environment rather than just the offline dataset. They then also directly tune the stepsize using the performance of the offline learned policy in the environment. Several papers (Kumar et al., 2019; Kostrikov et al., 2021) use the same methodology. Some works focus on designing algorithms that are robust to a specific hyperparameter (Cheng et al., 2022), or require only a small number of hyperparameters to be tuned (Fujimoto and Gu, 2021). These approaches, however, still require some hyperparameters to be tuned; they were tuned using the performance in the environment.
There are, however, a handful of works that use only the offline dataset to select hyperparameters. These can be categorized as papers focused on algorithms for hyperparameter selection in offline RL (Paine et al., 2020; Tang and Wiens, 2021; Yang et al., 2022), papers that provide theoretical algorithms in specialized settings (Farahmand and Szepesvári, 2011; Lee et al., 2022a), or papers that introduce offline RL algorithms and use sensible heuristics to select hyperparameters in their experiments (Yu et al., 2021; Kumar et al., 2021; Trabucco et al., 2021; Qi et al., 2022). For the first category, these algorithms are all based on OPE. For the second category, these algorithms are applicable only in specialized settings and require strong conditions to obtain the provided theoretical guarantees. For the last category, the heuristics are often tailored to the specific algorithm presented in these papers and do not come with performance guarantees. For example, Kumar et al. (2021) used statistics about the TD errors to select the number of training steps for CQL.
In general, there are as yet many open questions about the feasibility of hyperparameter selection in offline RL, both in theory and in practice. This is doubly true if we go beyond hyperparameter selection and consider selecting amongst different policy learning algorithms as well, each with their own hyperparameters. In this paper, we consider this more general problem, called offline policy selection (OPS), that can be used to select policy learning algorithms and their hyperparameters. Figure 2 shows the full offline RL pipeline with OPS.
Figure 2: The offline RL pipeline with OPS. In the policy training phase, $n$ algorithm-hyperparameter pairs are trained on an offline dataset to produce $n$ candidate policies. These $n$ policies, combined again with offline data (potentially a validation dataset) are run with an OPS algorithm to find a final policy amongst the candidate policies.
As mentioned above, OPE is a general approach to OPS, but we want to understand when, or even if, alternative approaches can outperform OPE for OPS. OPE might not be the ideal answer to OPS. One reason is that OPE requires a large number of samples to evaluate any given policy in the worst case (Wang et al., 2021). We might intuit that OPS may, at least in some cases, be easier than OPE, since OPS need only identify the best-performing policy, whereas OPE aims to estimate each policy’s value accurately. For example, Doroudi et al. (2017) designed a specialized importance sampling (IS) estimator to compare two policies, to remove a bias towards policies that result in shorter trajectories, improving on the standard IS estimator used in OPE. Therefore, the first question that needs to be answered is: Is OPS easier than OPE? Surprisingly, this basic question has not yet been explicitly answered.
Another key reason current OPE estimators do not resolve the OPS problem is because these OPE methods themselves have hyperparameters. For example, the MAGIC estimator (Thomas and Brunskill, 2016) and the clipped importance sampling (IS) estimator (Bottou et al., 2013) are sensitive to their hyperparameters. Even a variant of the IS estimators designed specifically for OPS (Yang et al., 2022) has hyperparameters. Other methods like FQE require learning a value function, and how to select the function class is still an open problem.
The natural alternative to OPE is to use Bellman errors (BE), but there is mixed evidence on whether such approaches are effective. Tang and Wiens (2021) empirically show that selecting the candidate value function with the smallest TD errors perform poorly because the value functions are often overestimated; they conclude OPE is necessary. Similarly, Paine et al. (2020) present positive experimental results using FQE—an OPE approach—for OPS but concluded that the use of TD errors was ineffective. Other works, however, have developed new OPS algorithms that rely on the use of BE. Zhang and Jiang (2021) propose a value-function selection algorithm called BVFT, which computes the (empirical) projected BE for each pair of candidate value functions. Lee et al. (2022b) provide a method for selecting the best function class from a nested set of function classes for Fitted Q-Iteration. These theoretically-sound methods, however, either rely on strong assumptions or are applicable only in specialized settings. It is unclear whether these methods are better than OPE.
Contributions.
In this work, we make contributions towards answering the above question. Our first contribution is to prove that OPS inherits the same hardness results as OPE, which has never been formally shown in the literature. We show that the sample complexity of the OPS problem is lower-bounded by the samples needed to perform OPE. We further show that an OPS algorithm that simply chooses the policy with the highest IS estimate achieves a nearly minimax sample complexity, which is exponential in the horizon. This result implies no OPS approach can avoid the exponential sample complexity in the worst case, and we must consider additional assumptions to enable sample efficient OPS.
We then investigate the utility of using BE for OPS. We propose a simple BE algorithm, called Identifiable BE selection (IBES), with a simple way to select its own hyperparameters. This is contrast to many OPE methods, even simpler approaches like FQE, for which it can be hard to select hyperparameters. We show that IBES can provide improved sample efficiency, but under stricter requirements on data coverage and on the candidate set compared to FQE. In an empirical study, we systematically compare different BE methods with varying sample sizes and show IBES is consistently more sample efficient across multiple environments.
From our empirical study, we also show the inherent difficulty of OPS in a larger scale experiment in Atari benchmark. Notably, we show that all the OPS methods suffer relatively high regret compared to using the tuned hyperparameters, in many cases performing even worse than random hyperparameter selection. Recall that the tuned hyperparameters are obtained by using performance in the real environment, which is not feasible in practice but nonetheless commonplace in offline RL experiments. This result in Atari highlights that many results in the literature report unrealistically favorable results for their offline algorithms, because they show results for these tuned hyperparameters rather than hyperparameters chosen using only the offline dataset.
2 Background
In RL, the agent-environment interaction can be formalized as a finite horizon finite Markov decision process (MDP) $M=(\mathcal{S},\mathcal{A},H,\nu,Q)$. $\mathcal{S}$ is a set of states with size $S=|\mathcal{S}|$, and $\mathcal{A}$ is a set of actions with size $A=|\mathcal{A}|$, $H\in{\mathbb{Z}}^{+}$ is the horizon, and $\nu\in\Delta(\mathcal{S})$ is the initial state distribution where $\Delta(\mathcal{S})$ is the set of probability distributions over $\mathcal{S}$. Without loss of generality, we assume that there is only one initial state $s_{0}$. The reward $R$ and next state $S^{\prime}$ are sampled from $Q$, that is, $(R,S^{\prime})\sim Q(\cdot|s,a)$. We assume the reward is bounded in $[0,r_{max}]$ almost surely so the total return of each episode is bounded in $[0,V_{max}]$ almost surely with $V_{max}=Hr_{max}$. The stochastic kernel $Q$ induces a transition probability $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, and a mean reward function $r(s,a)$ which gives the mean reward when taking action $a$ in state $s$.
A non-stationary policy is a sequence of memoryless policies $(\pi_{0},\dots,\pi_{H-1})$ where $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$. We assume that the set of states reachable at time step $h$, $\mathcal{S}_{h}\subset\mathcal{S}$, are disjoint, without loss of generality, because we can always define a new state space $\mathcal{S}^{\prime}=\mathcal{S}\times[H]$, where $[H]$ denotes the set $[H]:=\{0,1,\dots,H-1\}$. Then, it is sufficient to consider stationary policies $\pi:\mathcal{S}\to\Delta(\mathcal{A})$.
Given a policy $\pi$, for any $h\in[H]$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the value function and the action-value function as $v_{h}^{\pi}(s):={\mathbb{E}}^{\pi}[\sum_{t=h}^{H-1}r(S_{t},A_{t})|S_{h}\!=\!s]$ and $q_{h}^{\pi}(s,a):={\mathbb{E}}^{\pi}[\sum_{t=h}^{H-1}r(S_{t},A_{t})|S_{h}\!=\!s,A_{h}\!=\!a]$, respectively. The expectation is with respect to ${\mathbb{P}}^{\pi}$, which is the probability measure on the random element $(S_{0},A_{0},R_{0},\dots,R_{H-1})$ induced by the policy $\pi$. Similar to the stationary policy, we also use an action-value function $q$ to denote the sequence of action-value function $(q_{0},\dots,q_{H-1})$ since the set of states reachable at each time step are disjoint. That is, for all $h\in[H]$ and $s\in\mathcal{S}_{h}$, $q(s,\cdot)=q_{h}(s,\cdot)$.
The Bellman evaluation operator $\mathcal{T}^{\pi}$ is defined as
$(\mathcal{T}^{\pi}q)(s,a)=r(s,a)+\sum_{s^{\prime}\in\mathcal{S}}P(s,a,s^{\prime})\sum_{a^{\prime}\in\mathcal{A}}\pi(a^{\prime}|s^{\prime})q(s^{\prime},a^{\prime}),$
with $(\mathcal{T}^{\pi}q)(s,\cdot)=r(s,\cdot)$ for $s\in\mathcal{S}_{H-1}$. The Bellman optimality operator $(\mathcal{T}q)(s,a)$ is obtained if a greedy policy with respect to $q$ is used. We use $J(\pi)$ to denote the value of the policy $\pi$, that is, the expected return from the initial state $J(\pi)=v^{\pi}_{0}(s_{0})$. The optimal value function is defined by $v_{h}^{*}(s):=\sup_{\pi}v_{h}^{\pi}(s)$. A policy $\pi$ is optimal if $J(\pi)=v^{*}_{0}(s_{0})$.
In the offline setting, we are given a fixed set of transitions $D$ with samples drawn from a data distribution $\mu$. We consider the setting where the data is collected by a behavior policy $\pi_{b}$ since the data collection scheme is more practical (Xiao et al., 2022) and is used to collect data for benchmark datasets (Fu et al., 2020). We use $d^{\pi}_{h}$ to denote the data distribution at the horizon $h$ by following the policy $\pi$, that is, $d^{\pi}_{h}(s,a):={\mathbb{P}}^{\pi}(S_{h}=s,A_{h}=a)$, and $\mu_{h}(s,a):={\mathbb{P}}^{\pi_{b}}(S_{h}=s,A_{h}=a)$.
3 On the Sample Complexity of OPS
We consider the offline policy selection (OPS) problem and off-policy policy evaluation (OPE) problem. We follow a similar notation and formulation used in Xiao et al. (2022) to formally describe these problem settings. The OPS problem for a fixed number of episodes $n$ is given by the tuple $(\mathcal{S},\mathcal{A},H,\nu,n,{\mathcal{I}})$. ${\mathcal{I}}$ is a set of instances of the form $(M,d_{b},\Pi)$ where $M\in{\mathcal{M}}(\mathcal{S},\mathcal{A},H,\nu)$ specifies an MDP with state space $\mathcal{S}$, action space $\mathcal{A}$, horizon $H$ and the initial state distribution $\nu$, $d_{b}$ is a distribution over a trajectory $(S_{0},A_{0},R_{0},\dots,R_{H-1})$ by running the behavior policy $\pi_{b}$ on $M$, and $\Pi$ is a finite set of candidate policies. We consider the setting where $\Pi$ has a small size and does not depend on $S$, $A$ or $H$.
An OPS algorithm takes as input a batch of data $D$, which contains $n$ trajectories, and a set of candidate policies $\Pi$, and outputs a policy $\pi\in\Pi$. The aim is to find OPS algorithms that are guaranteed to return the best candidacy policy with high probability.
Definition 1 ($(\varepsilon,\delta)$-sound OPS algorithm).
Given $\varepsilon>0$ and $\delta\in(0,1)$, an OPS algorithm ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound on instance $(M,d_{b},\Pi)$ if
$\mathop{\rm Pr}\nolimits_{D\sim d_{b}}(J_{M}({\mathcal{L}}(D,\Pi))\geq J_{M}(\pi^{\dagger})-\varepsilon)\geq 1-\delta,$
where $\pi^{\dagger}$ is the best policy in $\Pi$. We say an OPS algorithm ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound on the problem $(\mathcal{S},\mathcal{A},H,\nu,n,{\mathcal{I}})$ if it is sound on any instance $(M,d_{b},\Pi)\in{\mathcal{I}}$.
Given a pair $(\varepsilon,\delta)$, the sample complexity of OPS is the smallest integer $n$ such that there exists a behavior policy $\pi_{b}$ and an OPS algorithm ${\mathcal{L}}$ such that ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound on the OPS problem $(\mathcal{S},\mathcal{A},H,\nu,n,{\mathcal{I}}(\pi_{b}))$ where ${\mathcal{I}}(\pi_{b})$ denotes the set of instances with data distribution $d_{b}$. That is, if the sample complexity is lower-bounded by a number $\mathrm{N}_{OPS}$, then, for any behavior policy $\pi_{b}$, there exists an MDP $M$ and a set of candidate policies $\Pi$ such that any $(\varepsilon,\delta)$-sound OPS algorithm on $(M,d_{b},\Pi)$ requires at least $\mathrm{N}_{OPS}$ episodes.
Similarly, the OPE problem for a fixed number of episodes $n$ is given by $(\mathcal{S},\mathcal{A},H,\nu,n,{\mathcal{I}})$. ${\mathcal{I}}$ is a set of instances of the form $(M,d_{b},\pi)$ where $M$ and $d_{b}$ are defined as above, and $\pi$ is a target policy. An OPE algorithm takes as input a batch of data $D$ and a target policy $\pi$, and outputs an estimate of the policy value. The goal is to estimate the true value accurately.
Definition 2 ($(\varepsilon,\delta)$-sound OPE algorithm).
Given $\varepsilon\in(0,V_{max}/2)$ and $\delta\in(0,1)$, an OPE algorithm ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound on instance $(M,d_{b},\pi)$ if
$\mathop{\rm Pr}\nolimits_{D\sim d_{b}}(|{\mathcal{L}}(D,\pi)-J_{M}(\pi)|\leq\varepsilon)\geq 1-\delta.$
We say an OPE algorithm ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound on the problem $(\mathcal{S},\mathcal{A},H,\nu,n,{\mathcal{I}})$ if it is sound on any instance $(M,d_{b},\pi)\in{\mathcal{I}}$.
Note that $\varepsilon$ should be less than $V_{max}/2$ otherwise the bound is trivial.
3.1 OPE as Subroutine for OPS
It is obvious that a sound OPE algorithm can be used for OPS, since we can run the sound OPE algorithm to evaluate each candidate policy and select the policy with the highest estimate. As a result, the sample complexity of OPS is upper-bounded by the sample complexity of OPE up to a logarithmic factor. For completeness, we state this formally in the theorem below.
Theorem 3 (Upper bound on sample complexity of OPS).
Given an MDP $M$, a data distribution $d_{b}$, and a set of policies $\Pi$, suppose that, for any pair $(\varepsilon,\delta)$, there exists an $(\varepsilon,\delta)$-sound OPE algorithm ${\mathcal{L}}$ on any OPE instance $I\in\{(M,d_{b},\pi):\pi\in\Pi\}$ with a sample size at most $O(\mathrm{N}_{OPE}(S,A,H,1/\varepsilon,1/\delta))$. Then there exists an $(\varepsilon,\delta)$-sound OPS algorithm for the OPS problem instance $(M,d_{b},\Pi)$ which requires at most $O(\mathrm{N}_{OPE}(S,A,H,2/\varepsilon,|\Pi|/\delta))$ episodes.
In terms of the sample complexity, we have an extra $\sqrt{\log{|\Pi|}/n}$ term for OPS due to the union bound. For hyperparameter selection in practice, the size of the candidate set is often much smaller than $n$, so this extra term is negligible. However, if the set is too large, complexity regularization (Bartlett et al., 2002) may need to be considered. In this paper, we only consider a finite candidate set.
3.2 OPS is not Easier than OPE
We have shown that OPS is sample efficient when OPE is sample efficient. However, it remains unclear whether OPS can be sample efficient when OPE is not. In the following theorem, we lower bound the sample complexity of OPS by the sample complexity of OPE. As a result, both OPS and OPE suffer from the same hardness result.
(a) Constructing $M^{\prime}$ and a candidate policy set
(b) Searching for the true policy value
Figure 3: Visual depiction of the reduction of OPE to OPS. Given a MDP $M$ and a target policy $\pi$, we can construct a new MDP $M^{\prime}$ and two candidate policies $\{\pi_{1},\pi_{2}\}$ for OPS, as shown in (a). The MDP construction was first mentioned in Wang et al. (2021). $\pi_{1}$ chooses $a_{1}$ in $s_{0}$ and is otherwise arbitrary, $\pi_{2}$ chooses $a_{2}$ and is otherwise identical to the target policy $\pi$. Figure (b) describes the search procedure to find the policy value by calling the OPS subroutine. When the OPS query returns $\pi_{1}$, we follow the green arrow. When the OPS query returns $\pi_{2}$, we follow the blue arrow. We can keep searching for the true policy value by setting $r$ for the OPS query, until the desired precision is reached.
Theorem 4 (Lower bound on sample complexity of OPS).
Suppose for any data distribution $d_{b}$ and any pair $(\varepsilon,\delta)$ with $\varepsilon\in(0,V_{max}/2)$ and $\delta\in(0,1)$, there exists an MDP $M$ and a policy $\pi$ such that any $(\varepsilon,\delta)$-sound OPE algorithm requires at least $\Omega(\mathrm{N}_{OPE}(S,A,H,1/\varepsilon,1/\delta))$ episodes. Then there exists an MDP $M^{\prime}$ with $S^{\prime}=S+2$, $H^{\prime}=H+1$, and a set of candidate policies such that for any pair $(\varepsilon,\delta)$ with $\varepsilon\in(0,V_{max}/3)$ and $\delta\in(0,1/m)$ where $m:=\lceil\log(V_{max}/\varepsilon)\rceil\geq 1$, any $(\varepsilon,\delta)$-sound OPS algorithm also requires at least $\Omega(\mathrm{N}_{OPE}(S,A,H,3/2\varepsilon,1/m\delta))$ episodes.
The proof sketch is to construct an OPE algorithm that queries OPS as a subroutine, as demonstrated in Figure 3. As a result, the sample complexity of OPS is lower bounded by the sample complexity of OPE. The proof can be found in Appendix A.
There exist several hardness results for OPE in tabular settings and with linear function approximation (Yin and Wang, 2021; Wang et al., 2021). Theorem 4 implies that the same hardness results hold for OPS since lower bounds for OPE are also lower bounds for OPS. We should not expect to have a sound OPS algorithm without additional assumptions to make OPE work. Theorem 4, however, does not imply that OPS and OPE are always equally hard. There are instances where OPS is easy but OPE is not. For example, when all policies in the candidate set all have the same value, any random policy selection is sound. However, OPE can still be difficult in such cases.
4 Minimax Sample Complexity for Finite Horizon Finite MDPs
In the previous section, we show that OPS and OPE problem are equally difficult in the worst case scenario. In this section, we show a lower bound on sample complexity of OPS for finite horizon finite MDPs, and that an OPE method (nearly) matches the lower bound.
We first present an exponential lower bound, using the lower bound construction from Xiao et al. (2022) and Theorem 4.
Corollary 5 (Lower bound on the sample complexity of OPS).
For any positive integers $S,A,H$ with $S>2H$ and a pair $(\varepsilon,\delta)$ with $0<\varepsilon\leq\sqrt{1/8}$, $\delta\in(0,1)$, any $(\varepsilon,\delta)$-sound OPS algorithm needs at least $\tilde{\Omega}(A^{H-1}/\varepsilon^{2})$ episodes.
We show that importance sampling (IS) achieves the lower bound. Recall the IS estimator (Rubinstein, 1981) is given by
$\hat{J}(\pi)=\frac{1}{n}\sum_{i=1}^{n}\prod_{h=0}^{H-1}\pi(A_{h}^{(i)}|S_{h}^{(i)})/\pi_{b}(A_{h}^{(i)}|S_{h}^{(i)})G_{i}$
where $G_{i}=\sum_{h=0}^{H-1}R_{h}^{(i)}$ and $n$ is the number of episodes in the dataset $D$.
Corollary 6 (Upper bound on the sample complexity of OPS).
Suppose the data collection policy is uniformly random, that is, $\pi_{b}(a|s)=1/A$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$, and $|G_{i}|\leq V_{max}$ almost surely. Then the selection algorithm ${\mathcal{L}}$ that selects the policy with the highest IS estimate is $(\varepsilon,\delta)$-sound with $O(A^{H}V_{max}\ln{(|\Pi|/\delta})/\varepsilon^{2})$ episodes.
Note that Wang et al. (2017) have shown that IS estimator achieved the minimax mean squared error for the OPE problem. Our result shows that IS also achieves a (nearly) minimax sample complexity for the OPS problem, which is different from their work.
These results suggest that IS achieves a nearly minimax optimal sample complexity for OPS up to a factor $A$ and logarithmic factors. There are other improved variants of IS, including per-decision IS and weighted IS (Precup et al., 2000; Sutton and Barto, 2018). None of these variants can help reduce sample complexity in the worst case because the lower bound in Corollary 5 holds for any OPS algorithm. This result suggests that we need to consider additional assumptions on the environment, the data distribution, or the candidate set to perform sample efficient OPS.
One direction is to consider when OPE can be sample efficient, since we can leverage Theorem 3 to inherit the sample complexity result from OPE to OPS. There is a wealth of literature on OPE, in particular, FQE and MIS (or DICE) methods have been shown to be effective for OPS empirically (Paine et al., 2020; Yang et al., 2022). FQE and MIS methods require a standard data coverage assumption, that is, we need data coverage for all the candidate policies. Data coverage is commonly measured by the concentration coefficient, first introduced in Munos (2007). [^1] There are other measures of data coverage (Yin and Wang, 2021; Xie et al., 2021). However, we focus on the concentration coefficient to clearly compare assumptions between different methods. The precise definition will be provided in the next section.
Besides data coverage, we also need additional assumptions on the function approximation. For FQE, we need a function class $\mathcal{F}$ that is closed under $\mathcal{T}^{\pi}$ for all $\pi\in\Pi$, that is, $\mathcal{T}^{\pi}q\in\mathcal{F}$ for any $q\in\mathcal{F}$. Assume $\mathcal{F}$ has finitely many elements (for simplicity only), the result from Duan et al. (2021) can be extended to show that the sample complexity of using FQE for OPS is $O(H^{4}\log{(|\mathcal{F}||\Pi|H/\delta)}/\varepsilon^{2})$, under both assumptions on data coverage and function class. There are corresponding results for the MIS (or DICE) methods (Uehara et al., 2020; Nachum et al., 2019). See Table 1 for a brief summary.
In summary, we showed an exponential lower bound on the sample complexity for OPS, and we need to consider additional assumptions such as data coverage and function approximation that enable sample efficient OPE.
5 Bellman Error Selection for OPS
In this section, we shift our focus to the use of Bellman error (BE) for OPS. There is mixed evidence on whether such approaches are effective compared to OPE. As a result, we first provide conditions when BE is useful for OPS, which allows us to clearly compare OPE and BE methods. We then propose a BE selection method that has a simple method to select its own hyperparameter, and proved the sample complexity of the method.
Suppose we are given a set of candidate value functions $\mathcal{Q}:=\{q_{1},\dots,q_{K}\}$ and let $\Pi=\{\pi_{1},\dots,\pi_{K}\}$ be the set of corresponding greedy policies, a common strategy is to select the action value function with the smallest BE (Farahmand and Szepesvári, 2011). In the finite horizon setting, we define the Bellman error with respect to $q_{i}$ as
$\mathcal{E}(q_{i}):=\frac{1}{H}\sum_{h=0}^{H-1}\|q_{i}-\mathcal{T}q_{i}\|_{2,\mu_{h}}^{2}$
where $\|q\|_{p,\mu_{h}}:=(\sum_{(s,a)\in\mathcal{S}_{h}\times\mathcal{A}}\mu_{h}(s,a)|q(s,a)|^{p})^{1/p}$. We define $(\varepsilon,\delta)$-sound BE selection in the following.
Definition 7 ($(\varepsilon,\delta)$-sound BE selection).
Given a set of candidate value functions $\mathcal{Q}$, $\varepsilon>0$ and $\delta\in(0,1)$, an BE selection algorithm ${\mathcal{L}}$, which takes $D,\mathcal{Q}$ as input and outputs $q\in\mathcal{Q}$, is $(\varepsilon,\delta)$-sound on $\mathcal{Q}$ if
$\mathcal{E}({\mathcal{L}}(D,\mathcal{Q}))\leq\min_{i=1,\dots,|\mathcal{Q}|}\mathcal{E}(q_{i})+\varepsilon$
with probability at least $1-\delta$.
5.1 When is BE Selection Useful for OPS?
In order to relate BE selection to OPS, we need data coverage for both the candidate policies $\Pi$ and an optimal policy (the first assumption in the following corollary). We present an error bound for OPS using BE selection.
Corollary 8 (Error amplification for OPS using BE selection).
Suppose
  1. 1.
    there exists a constant $C$ such that $\forall\pi\in\Pi\cup\{\pi^{*}\}$, $\underset{h\in[H]}{\max}\underset{s\in\mathcal{S}_{h},a\in\mathcal{A}_{h}}{\max}\frac{d_{h}^{\pi}(s,a)}{\mu_{h}(s,a)}\leq C$,
  2. 2.
    the suboptimality of the candidate set is small, that is, $\underset{q\in\mathcal{Q}}{\min}\ \mathcal{E}(q)\leq\varepsilon_{sub}$,
  3. 3.
    there exists an $(\varepsilon_{est},\delta)$-sound BE selection algorithm ${\mathcal{L}}$ on $\mathcal{Q}$,
then the OPS algorithm outputs the greedy policy with respect to ${\mathcal{L}}(D,\mathcal{Q})$ is $(2H\sqrt{C(\varepsilon_{sub}+\varepsilon_{est})},\delta)$-sound.
That is, even if we have a good BE selection algorithm with a small $\varepsilon_{est}$, the error for OPS is amplified by $C$ and $\varepsilon_{sub}$. Compared to OPE methods such as FQE, BE has an additional error $\varepsilon_{sub}$ which does not go to zero even when we can collect more samples for OPS. Only $\varepsilon_{est}$ goes to zero as $n$ goes to infinity, which will be discussed in the next subsection.
To see how poor the guarantee can be due to $\varepsilon_{sub}$, suppose we have two action values $q_{1}=100q^{*}$ and $q_{2}=q^{*}+\text{some random noise}$, then $q_{1}$ has a large Bellman error but $\pi_{1}$ is actually optimal. We would choose $q_{2}$ since it has a lower Bellman error, even when we can collect an infinite number of samples. To obtain a meaningful guarantee, we need to include another value function, for example, $q_{3}=q^{*}$ to make $\varepsilon_{sub}=0$. As we collect more samples, we can estimate the Bellman error more accurately and eventually choose $q_{3}$.
At the same time, as we get more offline data, then we might actually change our candidate set. If we have more samples for training offline algorithms and generating the candidate set, it is more likely that there is one action value function in the candidate set that is close to optimal. That is, for a fixed candidate set, $\varepsilon_{sub}$ does not get smaller as we collect more samples for OPS, but if we use those offline samples to train the candidates policies, then $\varepsilon_{sub}$ might get smaller.
5.2 A Sample-Efficient and Hyperparameter-Free BE Selection Method
In this subsection, we propose a sound BE selection method, which we call Identifiable BE Selection (IBES). In deterministic environments, the BE can be easily estimated using TD errors. Given a state-action value $q$ with $v_{q}(s):=\max_{a}q(s,a)$ the corresponding state value, the BE estimate is $\mathrm{TDE}(q):=\frac{1}{|D|}\sum_{(s,a,s^{\prime},r)\in D}(q(s,a)-r-v_{q}(s^{\prime}))^{2}$.
In stochastic environments, estimating BE typically involves an additional regression problem (Antos et al., 2008). Antos et al. (2008) propose an estimator for the BE by introducing an auxiliary function $g\in\mathcal{G}$:
^$\hat{\mathcal{E}}(q)$ $=\tfrac{1}{|D|}\!\!\sum_{(s,a,s^{\prime},r)\in D}\!\!(q(s,a)-r-v_{q}(s^{\prime}))^{2}-\min_{g\in\mathcal{G}}\tfrac{1}{|D|}\!\!\sum_{(s,a,s^{\prime},r)\in D}\!\!(g(s,a)-r-v_{q}(s^{\prime}))^{2}$
$=\max_{g\in\mathcal{G}}\left[\tfrac{1}{|D|}\!\!\sum_{(s,a,s^{\prime},r)\in D}\!\!(q(s,a)-r-v_{q}(s^{\prime}))^{2}-\tfrac{1}{|D|}\!\!\sum_{(s,a,s^{\prime},r)\in D}\!\!(g(s,a)-r-v_{q}(s^{\prime}))^{2}\right].$ (1)
The sample complexity of estimating the BE depends on the complexity of the function class $\mathcal{G}$, as shown in the theorem below.
Theorem 9.
Suppose all $q\in\mathcal{Q}$ and $g\in\mathcal{G}$ take value in $[0,V_{max}]$. Let $q^{\dagger}$ be the action value function with the smallest estimated Bellman error ^$\hat{\mathcal{E}}(q)$. Then with probability at least $1-\delta$, for some constant $c_{0}$ (ignoring approximation and optimization error),
$\mathcal{E}(q^{\dagger})\leq\min_{i=1,\dots,|\mathcal{Q}|}\mathcal{E}(q_{i})+c_{0}V_{max}^{2}\sqrt{\log(2|\Pi|H/\delta)/n}+c_{0}V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})$
where $\mathcal{R}_{n}^{\mu}(\mathcal{G})$ is the Rademacher complexity of the function class $\mathcal{G}$.
Assume $\mathcal{G}$ has finite elements (for simplicity only), the sample complexity of finding a good action value function is $O(H^{4}\log{(|\mathcal{G}||\Pi|H/\delta)}/\varepsilon_{est}^{2})$. Compared to deterministic environments, we have an additional term that depends on the size of the function class. The proof can be found in Appendix A.
We can rewrite the inner term in Equation (1) by the change of variable (Dai et al., 2018; Patterson et al., 2022). Let $t(r,s^{\prime}):=r+v_{q}(s^{\prime})$ be the target and $\delta(s,a,r,s^{\prime}):=t(r,s^{\prime})-q(s,a)$ be the TD error to simplify the equations. Consider a new auxiliary function $h(s,a)=g(s,a)-q(s,a)$, then the inner term in Equation (1) is
$\tfrac{1}{|D|}\sum_{(s,a,s^{\prime},r)\in D}\left[(q(s,a)-r-v_{q}(s^{\prime}))^{2}-(g(s,a)-r-v_{q}(s^{\prime}))^{2}\right]$
$=\tfrac{1}{|D|}\sum_{(s,a,s^{\prime},r)\in D}\left[(t(r,s^{\prime})-q(s,a))^{2}-(t(r,s^{\prime})-g(s,a))^{2}\right]$
$=\tfrac{1}{|D|}\sum_{(s,a,s^{\prime},r)\in D}\left[2h(s,a)\delta(s,a,r,s^{\prime})-h(s,a)^{2}\right].$
That is, we can also use an auxiliary function $h$ to predict $\mathcal{T}q-q$, instead of $\mathcal{T}q$, and use the auxiliary function to estimate the Bellman error. The benefit is that the Bellman errors are more likely to be predictable. Under the conditions for Corollary 8, if there is an action-value function $q\in\mathcal{Q}$ with a small BE, the Bellman errors are nearly zero everywhere and any reasonable function class are able to represent the solution. The remaining part is to find a function class that has low approximation error and a low statistical complexity.
Fortunately, we can perform model selection to choose the function approximation $\mathcal{G}$. This is because we are running regression with fixed targets in Eq (1), and model selection for regression is well-studied. For example, consider a finite set of potential function classes $\mathcal{G}_{1},\dots,\mathcal{G}_{M}$, we can use a holdout validation set to select the best function class and other hyperparameters. Therefore, we can choose a function class such that it has a low approximation error and a small complexity measure, which can potentially result in improved sample efficiency. We describe the full procedure of IBES in Algorithm 1.
Algorithm 1 Identifiable BE Selection (IBES) with holdout validation
Input: Candidate set $\mathcal{Q}$, training data $D$, validation data $D_{val}$, a set of function classes $\mathcal{G}_{1},\dots,\mathcal{G}_{M}$ (for model selection)
Let $\delta(s,a,r,s^{\prime}):=r+v_{q}(s^{\prime})-q(s,a)$
for $q\in\mathcal{Q}$ do
     for $m=1,\dots,M$ do
         Perform regression: $\hat{g}_{m}\leftarrow\min_{g\in\mathcal{G}_{m}}\frac{1}{|D|}\sum_{D}(g(s,a)-\delta(s,a,r,s^{\prime}))^{2}$
         Compute validation error: $l(\hat{g}_{m})\leftarrow\frac{1}{|D_{val}|}\sum_{D_{val}}(\hat{g}_{m}(s,a)-\delta(s,a,r,s^{\prime}))^{2}$
     Find the best function class: $k\leftarrow\operatorname*{arg\,min}_{m=1,\dots,M}l(\hat{g}_{m})$
     Estimate the Bellman error for $q$: $\mathrm{BE}(q)\leftarrow\frac{1}{|D|}\sum_{D}2\hat{g}_{k}(s,a)\delta(s,a,r,s^{\prime})-\hat{g}_{k}(s,a)^{2}$
Output: $q^{\dagger}\leftarrow\arg\min_{q\in\mathcal{Q}}\mathrm{BE}(q)$
5.3 Comparison to Existing Methods
There are other works consider selecting a value function that has the smallest BE or is the closest to the optimal value function. Farahmand and Szepesvári (2011) consider selecting a value function such that, with high probability, the output value functions has the smallest BE. They propose to fit a regression model $\tilde{q}_{i}$ to predict $\mathcal{T}q_{i}$ and bound the BE by $\|q_{i}-\tilde{q}_{i}\|_{2,\mu}^{2}+b_{i}$ where the first term can be viewed as the (empirical) projected Bellman error, and the second term $b_{i}$ is a high-probability upper bound on the excess risk of the regression model, which is assumed to be given. There is a concurrent work on using BE selection method for OPS (Zitovsky et al., 2023). They propose a method called Supervised Bellman Validation (SBV), which is essentially an empirical version of the method from Farahmand and Szepesvári (2011), without an additional upper bound on the excess risk. They did not provide any sample complexity guarantee. In our experiments, we find that our method outperforms SBV in terms of sample efficiency, likely because the auxiliary function is used to predict the Bellman error, instead of $\mathcal{T}q$.
Zhang and Jiang (2021) propose to use a (empirical) projected Bellman error, called BVFT loss, with piecewise constant function classes. Their selection algorithm chooses the value function with the smallest BVFT loss, assuming $q^{*}$ is in the candidate set (approximately) and a stronger data assumption is satisfied. Interestingly, this condition on having $q^{*}$ is essentially equivalent to our condition requiring small $\varepsilon_{sub}$, since $q^{*}$ has exactly zero BE. The algorithms, though, are quite different from our work. Their method is computationally expensive since it scales with $O(|\Pi|^{2})$ instead of $O(|\Pi|)$, making the method impractical when the candidate set is large. Our method requires a weaker data coverage assumption and the computation cost scales linearly with $O(|\Pi|)$. Note that the sample complexity of our methods depends on the function class that is used to perform the regression. BVFT can be viewed as using the piecewise constant function classes to fit the Bellman target, the sample complexity depends on the piecewise constant function classes, which is measured by the number of discretization bins $(V_{max}/\varepsilon_{dct})^{2}$ and $\varepsilon_{dct}$ is the discretization resolution.
5.4 Should We Use BE Selection or OPE?
Method Assumptions Sample Complexity
IS $\pi_{b}(a|s)>0$ if $\pi(a|s)>0$ for some $\pi\in\Pi$ $O(A^{H}H\ln{(|\Pi|/\delta})/\varepsilon^{2})$
FQE (1) Data coverage for $\Pi$ $O(H^{4}\log{(|\mathcal{F}||\Pi|H/\delta)}/\varepsilon^{2})$
(2) $\mathcal{F}$ is closed under $\mathcal{T}^{\pi}$ for all $\pi\in\Pi$
IBES (1) Data coverage for $\Pi\cup\{\pi^{*}\}$ $O(H^{4}\log{(|\mathcal{F}||\Pi|H/\delta)}/\varepsilon_{est}^{2})$
(2) Small $\varepsilon_{sub}$
(3) $\mathcal{G}$ realizes $(\mathcal{T}q-q)$ for $q\in\mathcal{Q}$
BVFT (1) Stronger data coverage Number of partitions
(2) $q^{*}\in\mathcal{Q}$ (which implies small $\varepsilon_{sub}$)
Table 1: A summary of OPS methods. These sample complexity results are derived in this paper, except for the sample complexity of BVFT given in Zhang and Jiang (2021). The first two methods are OPE methods and the last two methods use action value functions. The assumptions for function approximation can be relaxed to hold approximately.
Table 1 summarizes the comparison of BE and OPE methods. OPS using IBES and BVFT requires stronger data coverage assumptions since it needs coverage not only for the candidates policies, but also $\pi^{*}$. Moreover, the guarantee for OPS using IBES and BVFT can be poor due to the additional term $\varepsilon_{sub}$, which does not decrease as we collect more samples for OPS. We need at least one of the action-value functions to be close to the optimal action-value function. For OPE, we can perform OPS even if none of the candidate policies are close to optimal. Therefore, OPE is a more robust and reliable method for OPS.
On the other hand, if we satisfy this stronger data coverage condition and have small $\varepsilon_{sub}$, then IBES has several advantages. IBES can be much more sample efficient in deterministic environments, or even in stochastic environments by choosing an appropriate function approximation. More importantly, for IBES, we are not plagued by the issue of having hard-to-specify hyperparameters. This is critical for the offline setting, where we cannot test different hyperparameter choices in the environment.
6 Experimental Results
In this section, we empirically investigate the different BE methods as well as FQE for OPS. The goal in the experiments is to gain a better understanding of how IBES and FQE perform when we vary important factors such as data coverage, sample size, and candidate policies. We first compare different BE-based methods to show the advantage of our proposed method IBES, and investigate the differences between IBES and FQE in classic control environments where we vary the data coverage. Finally, we perform an experiment on the Atari offline benchmark dataset.
To evaluate the performance of OPS, we consider the normalized top-$k$ regret used in Zhang and Jiang (2021). Top-$k$ regret is the gap between the best policy within the top-$k$ policies, and the best policy among all candidate policies. We then normalized the regret by the difference between the best and the worst policy, so the normalized regret is between 00 and $1$. The normalized regret can be interpreted as the percentage of degradation if we use offline selection instead of online selection, since online selection with Monte Carlo evaluation can always achieve a regret of zero. OPS corresponds to $k=1$; for most results we use $k=1$, but include some results for $k>1$. All hyperparameters for OPS methods are selected using only the datasets, not by peaking at performance on the real environment; for more details, see Appendix B.
6.1 Comparison between BE-based Methods
In the first set of experiments, we compare different BE-based methods for OPS. We conduct experiments on two standard RL environments: Acrobot and Cartpole. We also include the stochastic versions of these environments with sticky actions (Machado et al., 2018), which we call Stochastic Acrobot and Stochastic Cartpole. We generate a set of candidate policy-value pairs by running CQL with different hyperparameters on a batch of data collected by a trained policy with random actions taken $40\%$ of the time, and generate datasets for OPS that provide good data coverage. We then use either FQE or IBES for OPS. We also included SBV (Zitovsky et al., 2023) and BVFT (Zhang and Jiang, 2021), [^2] We modified the BVFT implementation from the author of Zhang and Jiang (2021) (https://github.com/jasonzhang929/BVFT_empirical_experiments/). and a random selection baseline which selects a policy uniformly at random from the candidate set.
Figure 4: Comparison between BE methods. The figure shows the normalized top-$1$ regret with varying sample size, averaged over 10 runs with one standard error. IBES consistently achieves the lowest regret across environments.
Figure 5: Comparison to BE with a fixed number of hidden units. The figure shows the normalized top-$1$ regret with varying sample size, averaged over 10 runs with one standard error. IBES with model selection consistently achieves the lowest regret across environments.
We first compare different BE methods, and investigate the effect of using the Bellman error $\mathcal{T}q-q$ as target. We include a baseline that use $\mathcal{T}q$ as target, called IBES-TQ. Note that we also perform model selection for SBV and IBES-TQ, similar to IBES. In Figure 4, we can see all BE methods converge to the same performance as sample size gets larger, but IBES using the Bellman error as target is much more sample efficient than IBES-TQ and SBV across all environment. This is likely due to the fact that Bellman error is easier to predict so using a smaller neural network is sufficient and has a better sample efficiency. We also found BVFT performs similar to TDE (which are overlapping in Cartpole, Acrobot and Stochastic Cartpole), which is likely due to the fact that BVFT with small discretization is equivalent to TDE.
Now we show that the model selection procedure is important for IBES, by comparing to IBES with a fixed number of hidden units. In our experiment, we use a two layer neural network model as the function approximation, and perform model selection to find the number of hidden units from the set $\{32,64,128,256\}$ for IBES. In Figure 5, we include two baselines called IBES-32 and IBES-256, which has the hidden units of size 32 and 256 respectively. In general, IBES with model selection achieves regret less than or equal to IBES with a fixed number of hidden units across all environments. The performance of IBES with model selection and IBES-256 converges as the sample size gets larger. This result shows an improvement in terms of sample efficiency due to model selection. In stochastic environments, IBES with a small number of hidden units does not work well even with a large sample size. The results suggest that the ability to perform model selection to balance approximation and estimation error is important to improve sample efficiency while achieving low regret.
In summary, the experiment suggests the proposed BE method is more same efficient, due to the fact that the method predict the Bellman error, and the model selection procedure.
6.2 Comparison between FQE and IBES under Different Data Coverage
In the second set of experiments, we aim to understand of how IBES and FQE perform when we vary important factors such as data coverage and sample size. We design two different datasets for the experiments: (a) well-covered data is generated such that all candidate policies are well-covered, and (b) well-covered data with optimal trajectories includes more diverse trajectories collected by an $\varepsilon$-greedy optimal policy (which used in the previous experiment).
Figure 6: Comparison between IBES and FQE under different sample size and data distributions. The figure shows the normalized top-$1$ regret with varying sample size, averaged over 10 runs with one standard error.
Figure 6 shows the results for top-$1$ regret with varying numbers of episodes. We first focus on the asymptotic performance ($n\approx 10^{3}$) across different datasets. FQE performs very well with a small regret on well-covered and diverse data. IBES performs better with optimal trajectories, especially in Cartpole and Stochastic Acrobot. This result matches our theoretical result that IBES requires a stronger data coverage for an optimal policy. Moreover, IBES often performs better than FQE with a small sample size, suggesting that it could offer a slightly better sample efficiency than FQE.
Investigating the results deeper, we observed that when IBES does not perform well (Acrobot with optimal trajectories), it is often the case that one of the value functions has the smallest Bellman error but it is not far from optimal. When FQE does not perform well (Stochastic Cartpole with optimal trajectories), it is often the case that one of the candidate policies is highly overestimated. Therefore, we include a simple two-stage method that first uses FQE to select $k_{1}$ policies, then use IBES to find the top-$k_{2}$ policies amongst the $k_{1}$ selected policies. We set $k_{1}=10$ and $k_{2}=1$ in our experiment. The idea is similar to the two-stage method proposed in Tang and Wiens (2021). We label the two-stage method as FQE+IBES in the figure. We can see that it performs consistently well, better than either method alone. We hypothesize the explanation is that FQE usually performs very well for top-$k$ selection with a large $k$, even though top-$1$ regret might be bad. We then use IBES to select from a subset of reasonably good candidate policies, where the candidate with the smallest error is more likely to be optimal. Further investigation about combining multiple OPS methods might be a promising research direction.
6.3 Comparison between FQE and IBES on Atari Datasets
Finally, we conduct experiments on benchmark Atari datasets to show the hardness of OPS for offline RL, and compare IBES to FQE in a more practical setting. We use the offline data on Breakout and Seaquest from the DQN replay dataset [^3] https://research.google/resources/datasets/dqn-replay/, a commonly used benchmark in offline RL. We sample $1$ million transitions from different learning stages to promote data coverage. Unlike our previous experiment, the data coverage might be poor due to the absent of explicit exploration to cover all candidate policies. We use $50\%$ of the data to generate a set of candidate policy-value pairs by running CQL with different number of gradient steps and different regularization coefficients, as specified in Kumar et al. (2020). We use the other $50\%$ data to perform OPS using FQE or IBES. We generate two candidate sets for each environment. The early-learning candidate contains policies with gradient steps in $\{50k,\cdots,500k\}$ and final-learning candidate contains policies with gradient steps in $\{550k,\dots,1000k\}$. In this experiment, we use the CQL and FQE implementation from the d3rlpy library (Seno and Imai, 2022).
Figure 7: Regret improvement over the random baseline on the Atari dataset. We show the regret compared to the random baseline. A positive value means the method outperforms random selection, and a negative value means the method performs worse than random selection. We show the distribution of the regret improvement across 5 random seeds (each seed is a single point in the violin plots) and the regret of the random baseline under the x-axis labels. None of these methods can consistently outperform the random baseline, showing the hardness of OPS.
Figure 7 shows the top-$1$ regret. We can see that these none of the OPS methods can consistently outperform the random baseline, showing the hardness of OPS in a practical situation where we can not control the data coverage. For Breakout with early-learning candidate set, we found that IBES tends to pick the candidate value function with a small number of gradient steps. This is likely due to the fact that none of the candidate value functions are close to optimal (that is, large $\varepsilon_{sub}$) and the value function with a small number of training steps has a small magnitude and hence a small estimated Bellman error. IBES performs better with the candidate set containing policies in the final learning stage, where it is more likely to contain a value function that is close to optimal. This shows the limitation of IBES, as discussed in the previous section. We can also see that FQE is more robust to the choice candidate set since it does not require one of the value function being optimal, as long as all candidate set are covered by the data. However, it is more sensitive to the data used to run FQE, and hence high variance across different random seeds.
7 Related Work
In this section we provide a more comprehensive survey of prior work on model selection for RL. In the online setting, model selection has been studied extensively across contextual bandits (Foster et al., 2019) to RL (Lee et al., 2021). In the online setting, the goal is to select model classes while balancing exploration and exploitation to achieve low regret, which is very different from the offline setting where no exploration is performed.
In the offline setting, besides using OPE and BE selection, other work on model selection in RL is in other settings: selecting models and selecting amongst OPE estimators. Hallak et al. (2013) consider model selection for model-based RL algorithms with batch data. They focus on selecting the most suitable model that generates the observed data, based on the maximum likelihood framework. Su et al. (2020) consider adaptive estimation ion for OPE when the candidate estimators can be ordered with monotonically increasing biases and decreasing confidence intervals.
In offline RL pipelines, we often split the offline data into training data for generating multiple candidate policies and validation data for selecting the best candidate policy. Nie et al. (2022) highlight the utility of performing multiple random data splits for OPS. They do not study the hardness or sample complexity of this procedure.
To the best of our knowledge, there is no previous work on understanding the fundamental limits for OPS in RL. There is one related work in the batch contextual bandit setting, studying the selection of a linear model (Lee et al., 2022a). They provide a hardness result suggesting it is impossible to achieve an oracle inequality that balances the approximation error, the complexity of the function class, and data coverage. Our work considers the more general problem, selecting a policy from a set of policies, in the RL setting.
8 Conclusion
In this paper, we made contributions towards understanding when OPS is feasible for RL. One of our main results—that the sample complexity of OPS is lower-bounded by the sample complexity of OPE—is perhaps expected. However, to our knowledge, this has never been formally shown. This result implies that without conditions to make OPE feasible, we cannot do policy selection efficiently. Our second contribution is the proposed IBES algorithm. We provide a sample complexity analysis, and empirically show that it is more sample efficient than existing BE-based methods for OPS.
We expect active research topics will be (1) to identify suitable conditions on the policies, environments and data, to make OPS sample efficient, (2) to design offline RL algorithms that have sound methods to select their own hyperparameters, and (3) to investigate how to combine multiple methods for a better OPS algorithm. In offline RL, we cannot select hyperparameters by testing in the real-world, and instead are limited to using the offline data. OPS is arguably one of the most critical steps towards bringing RL into the real-world, and there is much more to understand.
Appendix A Technical Details
A.1 Proof of Theorem 3
See 3
Proof  The OPS algorithm ${\mathcal{L}}(D,\Pi)$ for a given $(\varepsilon,\delta)$ works as follows: we query an $(\varepsilon^{\prime},\delta^{\prime})$-sound OPE algorithm for each policy in $\Pi$ and select the policy with the highest estimated value. That is, ${\mathcal{L}}(D,\Pi)$ outputs the policy $\bar{\pi}:=\operatorname*{arg\,max}_{\pi\in\Pi}\hat{J}(\pi)$, where $\hat{J}(\pi)$ is the value estimate for policy $\pi$ by the $(\varepsilon^{\prime},\delta^{\prime})$-sound OPE algorithm with data $D$.
By definition of an $(\varepsilon^{\prime},\delta^{\prime})$-sound OPE algorithm we have
$\mathop{\rm Pr}\nolimits_{D\sim d_{b}}(|\hat{J}(\pi)-J(\pi)|\leq\varepsilon^{\prime})\geq 1-\delta^{\prime},\forall\pi\in\Pi.$
Applying the union bound, we have
$\mathop{\rm Pr}\nolimits_{D\sim d_{b}}(\forall{\pi\in\Pi},|\hat{J}(\pi)-J(\pi)|\leq\varepsilon^{\prime})$ $\geq 1-\delta^{\prime}|\Pi|.$
Let $\pi^{\dagger}$ denote the best policy in the candidate set $\Pi$, that is, $\pi^{\dagger}:=\operatorname*{arg\,max}_{\pi\in\Pi}J(\pi)$. With probability $1-\delta^{\prime}|\Pi|$, we have
$J(\bar{\pi})$ $\geq\hat{J}(\bar{\pi})-\epsilon^{\prime}\geq\hat{J}(\pi^{\dagger})-\varepsilon^{\prime}\geq J(\pi^{\dagger})-2\varepsilon^{\prime}.$
The second inequality follows from the definition of $\bar{\pi}$. Finally, by setting $\delta^{\prime}=\delta/|\Pi|$ and $\varepsilon^{\prime}=\varepsilon/2$, we get
$\mathop{\rm Pr}\nolimits_{D\sim d_{b}}(J(\bar{\pi})\geq J(\pi^{\dagger})-\varepsilon)\geq 1-\delta.$
That is, ${\mathcal{L}}$ is an $(\varepsilon,\delta)$-sound OPS algorithm, and it requires $O(\mathrm{N}_{OPE}(S,A,H,2/\varepsilon,|\Pi|/\delta))$ episodes.
A.2 Proof of Theorem 4
See 4
Proof  Our goal is to construct an $(\varepsilon,\delta)$-sound OPE algorithm with $\delta\in(0,1)$ and $\varepsilon\in[0,V_{max}/2]$. To evaluate any policy $\pi$ in $M$ with dataset $D$ sampled from $d_{b}$, we first construct a new MDP $M_{r}$ with two additional states: an initial state $s_{0}$ and a terminal state $s_{1}$. Taking $a_{1}$ at $s_{0}$ transitions to $s_{1}$ with reward $r$. Taking $a_{2}$ at $s_{0}$ transitions to the initial state in the original MDP $M$. See Figure 3 in the main paper for visualization.
Let $\Pi=\{\pi_{1},\pi_{2}\}$ be the candidate set in $M_{r}$ where $\pi_{1}(s_{0})=a_{1}$ and $\pi_{2}(s_{0})=a_{2}$ and $\pi_{2}(a|s)=\pi(a|s)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Since $\pi_{1}$ always transitions to $s_{1}$, it never transitions to states in MDP $M$. Therefore, $\pi_{1}$ can be arbitrary for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. We can add any number of transitions $(s_{0},a_{1},r,s)$ and $(s_{0},a_{2},0,s)$ in $D$ to construct the dataset $D_{r}$ with distribution $d_{b,r}$ arbitrarily.
Suppose we have an $(\varepsilon^{\prime},\delta^{\prime})$-sound OPS algorithm, where we set $\varepsilon^{\prime}=2\varepsilon/3$ and $\delta^{\prime}=\delta/m$ with $m:=\lceil\log(V_{max}/\varepsilon^{\prime})\rceil$. Note that if this assumption does not hold, then it directly implies that the sample complexity of OPS is larger than $\Omega(\mathrm{N}_{OPE}(S,A,H,1/\varepsilon,1/\delta))$. Our strategy will be to iteratively set the reward $r$ of MDP $M_{r}$ and run our sound OPS algorithm on $\Pi$ and using bisection search to estimate a precise interval for $J(\pi)$.
The process is as follows. By construction, our OPS algorithm will output either $\pi_{1}$, which has value $J_{M_{r}}(\pi_{1})=r$, or output $\pi_{2}$, which has value $J_{M_{r}}(\pi_{2})=J_{M}(\pi)$. That is, it has the same value as $\pi$ in the original MDP. Let us consider the following two cases. Let $\pi^{\dagger}$ be the best policy in $\Pi$ for MDP $M_{r}$.
Case 1: the OPS algorithm selects $\pi_{1}$. We know, by definition of a sound OPS algorithm, that
$\mathop{\rm Pr}\nolimits(J_{M_{r}}(\pi_{1})\geq J_{M_{r}}(\pi^{\dagger})-\varepsilon^{\prime})\geq 1-\delta^{\prime}$
$\implies$ $\mathop{\rm Pr}\nolimits(r\geq\text{max}(r,J_{M_{r}}(\pi_{2}))-\varepsilon^{\prime})\geq 1-\delta^{\prime}$
$\implies$ $\mathop{\rm Pr}\nolimits(J_{M_{r}}(\pi_{2})\leq r+\varepsilon^{\prime})\geq 1-\delta^{\prime}.$
Case 2: the OPS algorithm selects $\pi_{2}$.
$\mathop{\rm Pr}\nolimits(J_{M_{r}}(\pi_{2})\geq J_{M_{r}}(\pi^{\dagger})-\varepsilon^{\prime})\geq 1-\delta^{\prime}$
$\implies$ $\mathop{\rm Pr}\nolimits(J_{M_{r}}(\pi_{2})\geq\text{max}(r,J_{M_{r}}(\pi_{2}))-\varepsilon^{\prime})\geq 1-\delta^{\prime}$
$\implies$ $\mathop{\rm Pr}\nolimits(J_{M_{r}}(\pi_{2})\geq r-\varepsilon^{\prime})\geq 1-\delta^{\prime}.$
Given this information, we describe the iterative process by which we produce the estimate $\hat{J}(\pi)$. We first set $U=V_{max},L=0$ and $r=\frac{U+L}{2}$ and run the sound OPS algorithm with input $D_{r}$ of sample size $n_{r}$ and the candidate set $\Pi$. Then if the selected policy is $\pi_{1}$, then we conclude the desired event $J(\pi)\leq r+\varepsilon^{\prime}$ occurs with probability at least $1-\delta^{\prime}$, and set $U$ equal to $r$. If the selected policy is $\pi_{2}$, then we know the desired event $J(\pi)\geq r-\varepsilon^{\prime}$ occurs with probability at least $1-\delta^{\prime}$, and set $L$ equal to $r$. We can continue the bisection search until the accuracy is less than $\varepsilon^{\prime}$, that is, $U-L\leq\varepsilon^{\prime}$, and the output value estimate is $\hat{J}(\pi)=\frac{U+L}{2}$.
If all desired events at each call occur, then we conclude that $L-\varepsilon^{\prime}\leq J(\pi)\leq U+\varepsilon^{\prime}$ and thus $|J(\pi)-\hat{J}(\pi)|\leq\varepsilon$. The total number of OPS calls is at most $m$. Setting $\delta^{\prime}=\delta/m$ and applying a union bound, we can conclude that with probability at least $1-\delta$, $|J(\pi)-\hat{J}(\pi)|\leq\varepsilon$.
Finally, since any $(\varepsilon,\delta)$-sound OPE algorithm on the instance $(M,d_{b},\pi)$ needs at least $\Omega(\mathrm{N}_{OPE}(S,A,H,\varepsilon,1/\delta))$ samples, the $(\varepsilon^{\prime},\delta^{\prime})$-sound OPS algorithm also needs at least $\Omega(N_{OPE}(S,A,H,\varepsilon,1/\delta))$, or equivalently $\Omega(\mathrm{N}_{OPE}(S,A,H,3/(2\varepsilon^{\prime}),1/(m\delta^{\prime})))$ samples for at least one of the instances $(M_{r},d_{b,r},\Pi)$.
A.3 Proof of Corollary 6
Theorem A.1 (Exponential lower bound on the sample complexity of OPE).
For any positive integers $S,A,H$ with $S>2H$ and a pair $(\varepsilon,\delta)$ with $0<\varepsilon\leq\sqrt{1/8}$, $\delta\in(0,1)$, any $(\varepsilon,\delta)$-sound OPE algorithm needs at least $\Omega(A^{H}\ln{(1/\delta)}/\varepsilon^{2})$ episodes.
Proof  We provide a proof which uses the construction from Xiao et al. (2022). They provide the result for the offline RL problem with Gaussian rewards. Here we provide the result for OPE problem with Bernoulli rewards since we assume rewards are bounded to match Theorem 6.
Figure 8: Lower bound construction.
We can construct an MDP with $S$ states, $A$ actions and $2H$ states. See Figure 8 for the construction. Given any behavior policy $\pi_{b}$, let $a_{h}=\operatorname*{arg\,min}_{a}\pi_{b}(a|s_{h})$ be the action that leads to the next state $s_{h+1}$ from state $s_{h}$, and all other actions lead to an absorbing state $s_{h}^{o}$. Once we reach an absorbing state, the agent gets zero reward for all actions for the remainder of the episode. The only nonzero reward is in the last state $s_{H-1}$. Consider a target policy that chooses $a_{h}$ for state $s_{h}$ for all $h=0,\dots,H-1$, and two MDPs where the only difference between them is the reward distribution in $s_{H-1}$. MDP 1 has Bernoulli distribution with mean $1/2$ and MDP 2 has Bernoulli distribution with mean $1/2-2\varepsilon$. Let ${\mathbb{P}}_{1}$ denote the probability measure with respect to MDP 1 and ${\mathbb{P}}_{2}$ denote the probability measure with respect to MDP 2.
Let $\hat{r}$ denote the OPE estimate by an algorithm ${\mathcal{L}}$. Define an event $E=\{\hat{r}<\frac{1}{2}-\varepsilon\}$. Then ${\mathcal{L}}$ is not $(\varepsilon,\delta)$-sound if $({\mathbb{P}}_{1}(E)+{\mathbb{P}}_{2}(E^{c}))/2\geq\delta$. This is because ${\mathcal{L}}$ is not $(\varepsilon,\delta)$-sound if either ${\mathbb{P}}_{1}(\hat{r}<\frac{1}{2}-\varepsilon)\geq\delta$ or ${\mathbb{P}}_{2}(\hat{r}>\frac{1}{2}-\varepsilon)\geq\delta$.
Using the Bretagnolle–Huber inequality (see Theorem 14.2 of Lattimore and Szepesvári (2020)), we know
$\frac{{\mathbb{P}}_{1}(E)+{\mathbb{P}}_{2}(E^{c})}{2}\geq\frac{1}{4}\exp{(-D_{KL}({\mathbb{P}}_{1},{\mathbb{P}}_{2}))}.$
By the chain rule for KL-divergence and the fact that ${\mathbb{P}}_{1}$ and ${\mathbb{P}}_{2}$ only differ in the reward for $(s_{H-1},a_{H-1})$, we have
$D_{KL}({\mathbb{P}}_{1},{\mathbb{P}}_{2})$ $={\mathbb{E}}_{1}\left[\sum_{i=1}^{n}{\mathbb{I}}\{S_{H-1}^{(i)}=s_{H-1},A_{H-1}^{(i)}=a_{H-1}\}\left(\frac{1}{2}\log{(\frac{1/2}{1/2-\varepsilon})}+\frac{1}{2}\log{(\frac{1/2}{1/2+\varepsilon})}\right)\right]$
$=\sum_{i=1}^{n}{\mathbb{P}}_{1}(S_{H-1}^{(i)}=s_{H-1},A_{H-1}^{(i)}=a_{H-1})\left(-\frac{1}{2}\log{(1-4\varepsilon^{2})}\right)$
$\leq\frac{n8\varepsilon^{2}}{A^{H}}$
The last inequality follows from $-\log{(1-4\varepsilon^{2})}\leq 8\varepsilon^{2}$ if $24\varepsilon^{2}\leq 1/$ 2 (Krishnamurthy et al., 2016) and ${\mathbb{P}}_{1}(S_{H-1}^{(i)}=s_{H-1},A_{H-1}^{(i)}=a_{H-1})<1/A^{H}$ from the construction of the MDPs.
Finally,
$\frac{{\mathbb{P}}_{1}(E)+{\mathbb{P}}_{2}(E^{c})}{2}\geq\frac{1}{4}\exp{(-\frac{n8\varepsilon^{2}}{A^{H}})}$
which is larger than $\delta$ if $n\leq A^{H}\ln(1/4\delta)/8\varepsilon^{2}$. As a result, we need at least $\Omega(A^{H}\ln{(1/\delta)}/\varepsilon^{2})$ episodes.
See 6
Proof  Since the policy is uniform random, we know $|W_{i}G_{i}|<A^{H}V_{max}$ almost surely. Moreover, the importance sampling estimator is unbiased, that is, ${\mathbb{E}}[W_{i}G_{i}]=J(\pi)$. Using the Bernstein’s inequality, we can show that the IS estimator satisfies
$\mathop{\rm Pr}\nolimits\left(\left|\hat{J}(\pi_{k})-J(\pi_{k})\right|\leq\frac{2A^{H}V_{max}\ln{(2/\delta)}}{3n}+\sqrt{\frac{2\mathrm{Var}(W_{i}G_{i})\ln{(2/\delta)}}{n}}\right)\geq 1-\delta$
for one candidate policy $\pi_{k}$. Using the union bound over all candidate policies, we have
$\mathop{\rm Pr}\nolimits\left(\left|\hat{J}(\pi_{k})-J(\pi_{k})\right|\leq\frac{2A^{H}V_{max}\ln{(2|\Pi|/\delta)}}{3n}+\sqrt{\frac{2{\mathbb{V}}(W_{i}G_{i})\ln{(2|\Pi|/\delta)}}{n}},\forall k\right)\geq 1-\delta.$
That is,
$\mathop{\rm Pr}\nolimits\left(J({\mathcal{L}}(D,\Pi))\geq J(\pi^{\dagger})-\frac{4A^{H}V_{max}\ln{(2|\Pi|/\delta)}}{3n}+\sqrt{\frac{8{\mathbb{V}}(W_{i}G_{i})\ln{(2|\Pi|/\delta)}}{n}}\right)\geq 1-\delta.$
For the variance term,
${\mathbb{V}}(W_{i}G_{i})$ $={\mathbb{E}}[W_{i}^{2}G_{i}^{2}]-{\mathbb{E}}[W_{i}G_{i}]^{2}\leq{\mathbb{E}}[W_{i}^{2}G_{i}^{2}]\leq\sqrt{{\mathbb{E}}[W_{i}^{2}]{\mathbb{E}}[G_{i}^{2}]}\leq A^{H}V_{max}.$
The second inequality follows from the Cauchy-Schwarz inequality. Therefore, if $n>32A^{H}V_{max}\ln{(2|\Pi|/\delta)}/\varepsilon^{2}$, ${\mathcal{L}}$ is $(\varepsilon,\delta)$-sound.
A.4 Proof of Corollary 8
We first present the telescoping performance difference lemma, originally from Theorem 2 of Xie and Jiang (2020) for discounting setting, and Lemma 3.2 of Duan et al. (2021) for finite horizon setting. In the following results, we will write $q_{h}(s,a)$ with the horizon $h$ explicitly for clarity.
Lemma A.2 (Lemma 3.2 of Duan et al. (2021)).
Assume there exists a constant $C$ such that, for any $\pi\in\Pi\cup\{\pi^{*}\}$, $\max_{h}\max_{s\in\mathcal{S}_{h},a\in\mathcal{A}_{h}}\frac{d_{h}^{\pi}(s,a)}{\mu_{h}(s,a)}\leq C$. For any $q\in\mathcal{Q}$, let $\pi$ denotes the greedy policy with respect to $q=(q_{0},\dots,q_{H-1})$, then
$J(\pi)\geq J(\pi^{*})-2H\sqrt{C}\sqrt{\frac{1}{H}\sum_{h=0}^{H-1}\|q_{h}-\mathcal{T}q_{h+1}\|_{2,\mu_{h}}}\ .$
See 8
Proof  Since $\mathcal{E}(q_{k})=\frac{1}{H}\sum_{h}\|q_{k,h}-\mathcal{T}q_{k,h+1}\|_{\mu}^{2}\leq\varepsilon_{sub}+\varepsilon_{est}$ with probability at least $1-\delta$, Lemma A.2 implies that
$J(\pi_{k})$ $\geq J(\pi^{*})-2H\sqrt{C(\varepsilon_{sub}+\varepsilon_{est})}\geq J(\pi^{\dagger})-2H\sqrt{C(\varepsilon_{sub}+\varepsilon_{est})}$
where $\pi^{\dagger}$ is the best performing policy in $\Pi$.
By definition, the OPS algorithm is $(2H\sqrt{C(\varepsilon_{sub}+\varepsilon_{est})},\delta)$-sound for this instance.
A.5 Proof of Theorem 9
Assumption A.4 (Approximation error).
For any $h\in[H]$ and any candidate function $f\in\mathcal{Q}$ (here we use $f$ instead of $g$ to avoid confusion between $q$ and $g$), we assume the approximation error is bounded by $\varepsilon_{apx}$, that is,
$\inf_{g\in\mathcal{G}}\|g-\mathcal{T}f\|_{2,\mu_{h}}^{2}\leq\varepsilon_{apx}.$
Definition A.5 (Rademacher complexity).
Given a function class $\mathcal{F}$, let $X=\{x_{1},\dots,x_{n}\}$ denotes $n$ fixed data points at horizon $h$ following the distribution $\mu_{h}$, the empirical Rademacher complexity is defined as
$R_{X}(\mathcal{F})={\mathbb{E}}\left[\sup_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(x_{i})\mid X\right]$
where the expectation is with respect to the Rademacher random variables $\sigma_{i}$. The Rademacher complexity is defined as $R_{n}^{\mu_{h}}(\mathcal{F})={\mathbb{E}}[R_{X}(\mathcal{F})]$ where the expectation is with respect to the $n$ data points. Finally, to simply the notation, we define $R_{n}^{\mu}(\mathcal{F})=\max_{h\in[H]}R_{n}^{\mu_{h}}(\mathcal{F})$ as the maximum Rademacher complexity over all horizons.
Definition A.6.
Given an action value function $f$, define the state value function $v_{f}(s)=\arg\max_{a\in\mathcal{A}}f(s,a)$, and four loss functions:
$L_{D_{h}}(g,f)$ $:=\frac{1}{n}\sum_{(s,a,r,s^{\prime})\in D_{h}}(g(s,a)-r-v_{f}(s^{\prime}))^{2}$
$L_{D}(g,f)$ $:=\frac{1}{H}\sum_{h}\frac{1}{n}\sum_{(s,a,r,s^{\prime})\in D_{h}}(g(s,a)-r-v_{f}(s^{\prime}))^{2}$
$L_{\mu_{h}}(g,f)$ $:={\mathbb{E}}_{(s,a,r,s^{\prime})\sim\mu_{h}}[(g(s,a)-r-v_{f}(s^{\prime}))^{2}]$
$L_{\mu}(g,f)$ $:=\frac{1}{H}\sum_{h}{\mathbb{E}}_{(s,a,r,s^{\prime})\sim\mu}[(g(s,a)-r-v_{f}(s^{\prime}))^{2}].$
Theorem 9 (Bellman error selection in stochastic environments).
Suppose all $f\in\mathcal{Q}$ and $g\in\mathcal{G}$ take value in $[0,V_{max}]$. Let $k=\arg\min_{i}L_{D}(f_{i},f_{i})-L_{D}(\hat{g}_{i},f_{i})$ where $\hat{g}_{i}=\arg\min_{g\in\mathcal{G}}L_{D}(g,f_{i})$. Then the following holds with probability at least $1-\delta$,
$\mathcal{E}(q_{k})\leq\min_{i=1,\dots,|\mathcal{Q}|}\mathcal{E}(q_{i})+c_{0}\varepsilon_{apx}+c_{0}V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})+c_{0}V_{max}^{2}\sqrt{\frac{2\log(2|\Pi|H/\delta)}{n}}$
for some constant $c_{0}>0$.
Proof  Fix a horizon $h$ and a target function $f$. By concentration with Rademacher complexity (e.g., Lemma G.1 of Duan et al. (2021)), with probability at least $21-\delta/$ 2, we know, for any $g\in\mathcal{G}$,
$|L_{D_{h}}(g,f)-L_{\mu_{h}}(g,f)|\leq 2\mathcal{R}_{n}^{\mu_{h}}(L_{D_{h}}\circ\mathcal{G})+V_{max}^{2}\sqrt{\frac{\log(2/\delta)}{2n}}.$
We note that the loss function $L_{D_{h}}$ is $(2V_{max})$-Lipschitz in its first argument, that is,
$|L_{D_{h}}(g,f)(s,a,r,s^{\prime})-L_{D_{h}}(g^{\prime},f)(s,a,r,s^{\prime})|$
$=|(g(s,a)-r-V_{f}(s^{\prime}))^{2}-(g^{\prime}(s,a)-r-V_{f}(s^{\prime}))^{2}|$
$=|(g(s,a)-g^{\prime}(s,a))(g(s,a)+g^{\prime}(s,a)-2r-2V_{f}(s^{\prime}))|$
$\leq|g(s,a)-g^{\prime}(s,a)||g(s,a)+g^{\prime}(s,a)-2r-2V_{f}(s^{\prime})|$
$\leq|g(s,a)-g^{\prime}(s,a)|2V_{max}.$
Therefore,
$|L_{D_{h}}(g,f)-L_{\mu_{h}}(g,f)|\leq 4V_{max}\mathcal{R}_{n}^{\mu_{h}}(\mathcal{G})+V_{max}^{2}\sqrt{\frac{\log(2/\delta)}{2n}}.$
By Hoeffding’s inequality, we also know, with probability at least $21-\delta/$ 2,
$|L_{D_{h}}(f,f)-L_{\mu_{h}}(f,f)|\leq V_{max}^{2}\sqrt{\frac{\log(2/\delta)}{2n}}.$
By the union bound over all $h\in[H]$, we have, with probability at least $1-\delta$,
$|L_{D}(g,f)-L_{\mu}(g,f)|\leq 4V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})+V_{max}^{2}\sqrt{\frac{\log(2H/\delta)}{2n}},\forall g\in\mathcal{G}\text{, and}$
$|L_{D}(f,f)-L_{\mu}(f,f)|\leq V_{max}^{2}\sqrt{\frac{\log(2H/\delta)}{2n}}.$
Recall that we define $\hat{g}$ as the empirical minimizer, that is, $\hat{g}=\arg\min_{g\in\mathcal{G}}L_{D}(g,f)$, and let $g^{\dagger}$ be the population minimizer, that is, $g^{\dagger}=\arg\min_{g\in\mathcal{G}}L_{\mu}(g,f)$. It follows that with probability at least $1-\delta$,
$|L_{D}(f,f)-L_{D}(\hat{g},f)-(L_{\mu}(f,f)-L_{\mu}(\mathcal{T}f,f))|$
$\leq|L_{D}(f,f)-L_{\mu}(f,f)|+|L_{D}(\hat{g},f)-L_{\mu}(\mathcal{T}f,f)|$
$\leq|L_{D}(f,f)-L_{\mu}(f,f)|+|\underbrace{L_{D}(\hat{g},f)-L_{D}(g^{\dagger},f)}_{\leq 0}|+|L_{D}(g^{\dagger},f)-L_{\mu}(g^{\dagger},f)|+$
$\ \ \ \ \underbrace{|L_{\mu}(g^{\dagger},f)-L_{\mu}(\mathcal{T}f,f)|}_{=\|g^{\dagger}-\mathcal{T}f\|_{2,\mu}^{2}}$
$\leq|L_{D}(f,f)-L_{\mu}(f,f)|+|L_{D}(g^{\dagger},f)-L_{\mu}(g^{\dagger},f)|+\varepsilon_{apx}$
$\leq\varepsilon_{apx}+4V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})+V_{max}^{2}\sqrt{\frac{2\log(2H/\delta)}{n}}.$
Note that $L_{\mu}(f,f)-L_{\mu}(\mathcal{T}f,f)=\mathcal{E}(f)$. Now apply the union bound for all $f$ in the candidate set by setting $\delta=\delta/|\Pi|$, then the following holds for all $f$
$|L_{D}(f,f)-L_{D}(\hat{g},f)-\mathcal{E}(f)|$ $\leq\varepsilon_{apx}+4V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})+V_{max}^{2}\sqrt{\frac{2\log(2|\Pi|H/\delta)}{n}}.$
Let $k=\arg\min_{i}L_{D}(f_{i},f_{i})-L_{D}(\hat{g}_{i},f_{i})$, the following holds with probability at least $1-\delta$,
$\mathcal{E}(f)\leq\min_{i=1,\dots,|\mathcal{Q}|}\mathcal{E}(f_{i})+c_{0}\varepsilon_{apx}+c_{0}V_{max}\mathcal{R}_{n}^{\mu}(\mathcal{G})+c_{0}V_{max}^{2}\sqrt{\frac{2\log(2|\Pi|H/\delta)}{n}}$
for some constant $c_{0}>0$.
Assume $\mathcal{F}$ has finite elements, and $\varepsilon_{apx}=0$, then we need a sample size of $n=O(H^{4}\log{(|\mathcal{F}||\Pi|H/\delta)}/\varepsilon_{est}^{2})$.
A.6 Sample Complexity of FQE for OPS
Consider a function class $\mathcal{F}=\mathcal{F}_{1}\times\dots\times\mathcal{F}_{H-1}$ such that $\mathcal{F}$ is closed under $\mathcal{T}^{\pi}$ for all $\pi\in\Pi$ and $|\mathcal{F}_{h}|$ is finite for all $h$. Assume $r_{max}=1$ for simplicity. Again, we will write $q_{h}(s,a)$ with the horizon $h$ explicitly for clarity. Given a policy $\pi\in\Pi$, we can show that
$\|q^{\pi}_{0}-q_{0}\|_{1,d^{\pi}_{0}}$
$=\sum_{a}\pi(a|s_{0})|q^{\pi}_{0}(s_{0},a)-q_{0}(s_{0},a)|$
$=\sum_{a}\pi(a|s_{0})|(\mathcal{T}^{\pi}q^{\pi}_{1})(s_{0},a)-(\mathcal{T}^{\pi}q_{1})(s_{0},a)+(\mathcal{T}^{\pi}q_{1})(s_{0},a)-q_{0}(s_{0},a)|$
$\leq\sum_{a,s^{\prime},a^{\prime}}\pi(a|s_{0})p(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})|q^{\pi}_{1}(s,a)-q_{1}(s,a)|+\sum_{a}\pi(a|s_{0})|(\mathcal{T}^{\pi}q_{1})(s_{0},a)-q_{0}(s_{0},a)|$
$=\|q^{\pi}_{1}-q_{1}\|_{1,d^{\pi}_{1}}+\|\mathcal{T}^{\pi}q_{1}-q_{0}\|_{1,d^{\pi}_{0}}.$
Apply the same inequality recursively, we have
$\|q^{\pi}_{0}-q_{0}\|_{1,d^{\pi}_{0}}$ $\leq\sum_{h=0}^{H-1}\|\mathcal{T}^{\pi}q_{h+1}-q_{h}\|_{1,d^{\pi}_{h}}$
$\leq H\sqrt{\frac{1}{H}\sum_{h=0}^{H-1}\|\mathcal{T}^{\pi}q_{h+1}-q_{h}\|_{2,d^{\pi}_{h}}^{2}}$
$\leq H\sqrt{C\frac{1}{H}\sum_{h=0}^{H-1}\|\mathcal{T}^{\pi}q_{h+1}-q_{h}\|_{2,\mu_{h}}^{2}}\ .$
The second inequality follows from the Cauchy-Schwarz inequality, and the last inequality follows from data coverage assumption.
Theorem 5.3 and Proposition 6.1 and of Duan et al. (2021) imply that for some constant $c_{0}$, with probability at least $1-\delta$,
$\left|J(\pi)-\sum_{a\in\mathcal{A}}\pi(a|s_{0})q_{0}(s_{0},a)\right|$ $\leq c_{0}H\sqrt{CH\left(\sum_{h}\frac{H\log{|\mathcal{F}_{h}|}}{n}+H^{2}\frac{\log{(H/\delta)}}{n}\right)}$
$\leq c_{0}H\sqrt{CH^{2}\frac{\log{(|\mathcal{F}|H/\delta)}}{n}}.$
Apply the union bound, we know the following holds for all $\pi\in\Pi$ with probability at least $1-\delta$
$\left|J(\pi)-\sum_{a\in\mathcal{A}}\pi(a|s_{0})q_{0}(s_{0},a)\right|$ $\leq c_{0}H\sqrt{CH^{2}\frac{\log{(|\mathcal{F}||\Pi|H/\delta)}}{n}}.$
To get an accuracy of $\varepsilon/2$, we need $n\geq c_{1}H^{4}C\log{(|\mathcal{F}||\Pi|H/\delta)}/\varepsilon^{2}$ for some constant $c_{1}$.
Appendix B Experimental Details
In this section, we provide the experimental details for classic RL environments and Atari environments.
B.1 FQE Implementation
Since it is unclear how to perform model selection for FQE, we fix the function approximation as a two layer neural network model with hidden size $256$ for classic RL experiment, and as a convolutional neural network followed by one layer neural network for atari experiments.
It is known that FQE can diverge, due to the fact that it combines off-policy learning with bootstrapping and function approximation, known as the deadly triad (Sutton and Barto, 2018). If one of the candidate policies is not well-covered, then the FQE estimate may overestimate the value of the uncovered policy (or even diverge to a very large value) and resulting in poor OPS. To circumvent the issue of uncovered policies, we need assign low value estimates for uncovered policies. In our FQE implementation, we assign low value estimates to policies for which FQE diverges so the OPS algorithm would not choose these policies. We provide a pseudocode for OPS using FQE in Algorithm 2. In our experiment, we set $U=V_{max}+100$.
Algorithm 2 OPS using FQE
Input: Candidate set $\Pi$, training data $D$, function class $\mathcal{F}$, threshold $U$
for $\pi\in\Pi$ do
     Initialize $q_{H-1}=0$
     for $h=H-2,\dots,0$, $q_{h}\leftarrow\arg\min_{f\in\mathcal{F}}\hat{l}_{h}(f,q_{h+1})$ where
$\hat{l}_{h}(f,q_{h+1}):=\frac{1}{|D_{h}|}\sum_{(s,a,r,s^{\prime})\in D_{h}}(f(s,a)-r-q_{h+1}(s^{\prime},\pi(s^{\prime})))^{2}$
     Estimate the policy value $\hat{J}(\pi)\leftarrow{\mathbb{E}}_{a\sim\pi(\cdot|s_{0})}[q_{0}(s_{0},a)]$
     if $\hat{J}(\pi)>U$ then
         $\hat{J}(\pi)\leftarrow-\infty$
Output: $\pi^{\dagger}\leftarrow\arg\max_{\pi\in\Pi}\hat{J}(\pi)$
B.2 Classic RL Experiments
Stochastic environments.
We implement stochastic environments by sticky actions. That is, when the agent selects an action to the environment, the action might be repeated with probability $25\%$, up to a maximum of $4$ repeats.
Generating candidate policies.
To generate a set of candidate policies, we run CQL with different hyperparameter configurations on a batch of data with $300$ episodes collected with an $\varepsilon$-greedy policy with respect to a trained policy where $\varepsilon=0.4$. The hyperparameter configuration includes:
  • Learning rate $\in\{0.001,0.0003,0.0001\}$
  • Network hidden layer size $\in\{128,256,512\}$
  • Regularization coefficient $\in\{1.0,0.1,0.01,0.001,0.0\}$
  • Iterations of CQL $\in\{100,200\}$
As a result, we have $90$ candidate policies.
Generating data for OPS.
To generate data for offline policy selection, we use three different data distributions: (a) a data distribution collected by running the mixture of all candidate policies. As a result, the data distribution covers all candidate policies well; and (b) a data distribution collected by running the mixture of all candidate policies and an $\varepsilon$-greedy optimal policy that provides more diverse trajectories than (a).
Randomness across multiple runs.
To perform experiments with multiple runs, we fix the offline data and the candidate policies and only resample the offline data for OPS. This better reflects the theoretical result that the randomness is from resampling the data for an OPS algorithm. In our experiments, we use $10$ runs and report the average regret with one standard error. Since the variability across runs is not large, we find using $10$ runs is enough.
Random selection baseline.
We include a random selection baseline that randomly chooses $k$ policies. Since the random selection algorithm has very high variance, we estimate the expected regret of random selection by repeating the random selection $10000$ times, and report the average regret.
BVFT.
BVFT has a hyperparameter: the discretization resolution. We follow the method described in the original paper to search for the best resolution from a set of predefined values. Note that in the authors’ implementation, they use different sets for different environments.
B.3 Atari Experiments
Generating candidate policies.
To generate a set of candidate policies, we run CQL with the hyperparameters used in the original paper:
  • Regularization coefficient $\in\{0.5,4.0,5.0\}$
  • Number of gradient steps $\in\{50k,100k,150k,\dots,1000k\}$
Randomness across multiple runs.
Similar to the classic RL experiments, we fix the candidate policies and only resample the offline data for OPS. In our experiments, we use $5$ different DQN replay dataset and report the average regret with one standard error.
B.4 Compute
For Atari experiments, we used NVIDIA Tesla V100 GPUs. Each run for generating candidate policies, running FQE, or running IBES took less than 6 hours. For classic RL experiments, we used only CPUs. Each run for generating candidate policies, running FQE, or running IBES took less than 3 hours.